### v3

# A Survey on Adversarial Contention Resolution

Abstract.
Contention resolution addresses the challenge of coordinating access by multiple processes to a shared resource such as memory, disk storage, or a communication channel. Originally spurred by challenges in database systems and bus networks, contention resolution has endured as an important abstraction for resource sharing, despite decades of technological change. Here, we survey the literature on resolving worst-case contention, where the number of processes and the time at which each process may start seeking access to the resource is dictated by an adversary. We also highlight the evolution of contention resolution, where new concerns—such as security, quality of service, and energy efficiency—are motivated by modern systems. These efforts have yielded insights into the limits of randomized and deterministic approaches, as well as the impact of different model assumptions such as global clock synchronization, knowledge of the number of processors, feedback from access attempts, and attacks on the availability of the shared resource.
survey, contention resolution, backoff, wakeup problem, selection problem, multiple access channel, broadcast, algorithms, stability, conflict resolution, Ethernet, makespan, queuing theory, adversarial queuing theory
conference: Full Conference Name; April 1, 2025; Locationbooktitle: Proceedings of the Full Conference Nameccs: Theory of computation Distributed algorithmsccs: Networks Network algorithmsccs: Networks Network protocolsccs: Security and privacy Network security
1. Introduction
  • “I saw two New Yorkers, complete strangers, sharing a cab. One guy took the tires and the radio; the other guy took the engine.”    – David Letterman
How do we share a resource? This is a common problem, not only in large cities such as New York City, but also in a wide range of computer systems such as: wireless clients in WiFi networks (bianchi2000performance, ) and cellular networks (RamaiyanDo14, ); processes in a shared-memory architecture (Ben-DavidB17, ); nodes deciding on membership in a maximal independent set (daum2013maximal, ); and concurrency control for databases (mitra:probabilistic, ; mitra:non-exlusive, ). These are examples where system participants must share a resource, such as memory, disk storage, a data structure, or a wired or wireless communication channel. Contention resolution refers to this challenge of coordinating access to a common resource.
The origins of contention resolution lie with initial research on databases and old Ethernet networks (kurose:computer, ). However, the problem has grown in importance with the adoption of wireless communication, where the resource under contention is a broadcast channel. This began with the ALOHA protocol (abramson1985development, ), which initiated an effort in the research community to understand the performance of ALOHA, particularly when packet arrival times are dictated by a stochastic process (for examples, see (ALOHAnet, ; tsybakov1979ergodicity, ; ferguson1975control, ; fayolle1977stability, ; abramson1977throughput, ; rosenkrantz1983instability, ; kelly1987number, ; kelly:decentralized, )). A foundational algorithm from this time period is randomized binary exponential backoff (BEB) (MetcalfeBo76, ), which continues to play an important role in wireless communication. Subsequently, a number of variants were designed and analyzed, such as polynomial backoff (haastad1996analysis, ). The core ideas used by these older algorithms of multiplicatively increasing and decreasing the sending probability (i.e., “backing on” and “backing off”) remain key building blocks for many recent results.
While models with stochastic arrivals received significant attention from the theory community, empirical studies emerged to suggest that traffic often exhibits bursty behavior (martin2023self, ; Teymori2005, ; yu:study, ), where a large batch of packets arrive simultaneously. Thus, researchers considered a setting where a batch of packets arrive together and must be successfully transmitted on the shared channel before the next batch arrives, and so on. While this burstiness is still somewhat well-behaved—since batches never overlap—it posed new challenges over the stochastic setting. A surprising result by Bender et al. (bender:adversarial, ) demonstrated that BEB performs sub-optimally, requiring $\Omega(n\log n)$ slots to complete a single batch of $n$ packets. This spurred interest in better algorithms for batched arrivals, and we discuss these results in Section 2.
Over the past decade and a half, attention has turned to the case where the arrival times of packets are far-less constrained. One common approach has been to view this problem through the lens of adversarial queuing theory (borodin2001adversarial, ), where results are parameterized by traffic characteristics such as the arrival rate and the degree of burstiness. Another major branch in the literature addresses worst-case arrival times. Arguably, deterministic approaches have been influenced by the pioneering result by Komlos and Greenberg (komlos:asymptotically, ), leading to new solutions for worst-case arrival times. A prime example is the work by De Marco and Kowalski (doi:10.1137/140982763, ), and there has been a long line of results in this area (e.g., (DBLP:conf/icdcs/MarcoKS19, ; Chlebus:2016:SWM:2882263.2882514, ; DEMARCO20171, ; de2013contention, )). In parallel, many different randomized approaches have been explored. A recent influential paper by Chang et al. (ChangJP19, ) shows how the classic ingredients of backing off/on can yield an elegant solution to worst-case arrivals. The literature in these areas is explored in Sections 3.1, 3.2 and 3.3.
More recently, the problem has evolved further to capture the challenges motivated by modern systems. For example, energy efficiency is important to many ad-hoc and mobile network settings, where devices are typically battery-powered (SINHA201714, ; kenny:performance, ; polastre:telos, ). Another concern is malicious interference, which amounts to a denial-of-service attack (pirayesh2022jamming, ; arjoune2020smart, ). Early work in this area by Awerbuch et al. (awerbuch:jamming, ), leading to subsequent results by Richa et al. (richa:jamming2, ; DBLP:journals/dc/RichaSSZ13, ; richa:jamming3, ; DBLP:journals/ton/RichaSSZ13, ; richa:jamming4, ) and Ogierman et al. (ogierman:competitive, ; ogierman2018sade, ), showed how such interference can be tolerated, and attack-resistant contention resolution has since become an active area of research. These issues and others discussed in Sections  3.3 and 4.
Scope and Goals. In this survey, we focus primarily on contention resolution in the presence of an adversary that (1) controls when participants can begin executing their contention resolution protocol, and (2) may deny access to the shared resource for periods of time; we refer to this as adversarial contention resolution. This stands in contrast to the setting where arrival times are governed by a stochastic process, which is an area of research that has been previously surveyed by Chlebus (chlebus2001randomized, ).
What motivates our focus on adversarial contention resolution? From a theoretical standpoint, it is interesting to explore what is possible in such a challenging setting, but there are also practical considerations. Network traffic does not always align with stochastic models; specifically, bursty traffic fails to be captured by well-behaved (e.g., Poisson- or Bernoulli-distributed) arrival times. Additionally, the mobility of many modern computing devices, along with connectivity challenges in wireless systems, can lead to situations where the contention for a shared resource changes rapidly over time. Such challenges are addressed by algorithms that guarantee performance results under the problem of adversarial contention resolution.
Given this motivation, our overall goal in this survey paper is to summarize the many models and performance metrics present in the literature on adversarial contention resolution, along with discussing their corresponding theoretical results.
1.1. Definitions, Terminology, Models, Metrics, and Algorithm Design
Contention resolution has gone by other names in the literature, such as $k$-selection (FernandezAntaMoMu13, ), all-broadcast (chlebus2001randomized, ), and conflict resolution (komlos:asymptotically, ). To provide a unified presentation, we review terminology that is common to the literature, elaborate on typical model assumptions and performance metrics, and highlight those aspects used to organize the literature.
1.1.1. Multiple Access Channel.
Throughout, we consider a time-slotted system with $n$ $\geq 2$ system participants, each seeking access to a shared resource. In each slot, a participant may choose to access the shared resource and, if it is the only participant to do so, this access attempt is successful. Else, if two or more participants make an attempt in the same slot, the result is a collision and no participant succeeds in this slot. Finally, if no participant makes an access attempt, then the corresponding slot is empty.
In practice, the resource itself may be local to a single machine; for example, corresponding to the critical section of a program, or a common disk shared by multiple processes. This extends to networked settings where participants access shared memory or separate machines via remote procedure calls.
In the literature, the resource is frequently treated as a shared communication channel, and this is the convention we use throughout this survey. Such a channel is typically referred to as a multiple access channel, which is a popular model under which several fundamental problems have been examined, such as mutual exclusion (bienkowski2010dynamic, ; czyzowicz2009consensus, ), leader election (chang:exponential-jacm, ; chang:exponential, ; nakano2002, ), consensus (czyzowicz2009consensus, ), routing (chlebus2019energy, ), broadcast problem (Marco08, ; Marco10, ; ChlebusGGPR00, ; alon1991lower, ; ClementiMS03, ; DBLP:conf/podc/Kowalski05, ; chrobak2002fast, ; piotr:explicit, ; chlebus2005almost, ; kowalski2005time, ), and well-known random access protocols (kurose:computer, ). Here, for any given slot, accessing the resource corresponds to a participant listening (equivalently, receiving) or sending (equivalently, transmitting) a packet on the multiple access channel. Unless stated otherwise, a packet can be transmitted within a single slot. When a packet is sent in a slot, we say this packet succeeds if no other packet is sent in the same slot; otherwise, the packet fails.
For ease of exposition, we will often slightly abuse terminology by speaking of packets themselves taking action (e.g., sending themselves on the channel, listening on the channel). This “packet-centric” presentation allows us to avoid referring to the corresponding “participants”/“devices”/“nodes” that actually performs these actions.
1.1.2. Models Assumptions.
We discuss model assumptions that play an important role in establishing asymptotic upper and lower bounds with respect to the metrics discussed above.
Packet Arrivals. An important aspect is the timing by which new packets are generated (i.e., arrive) in the system. A simple case is where each participant has a single packet, and all participants start executing a contention-resolution algorithm at the same point in time. Equivalently, from a packet-centric view, all packets arrive simultaneously as a single batch; this is often referred to as the static case in the literature (see (DBLP:conf/podc/Kowalski05, )).
In contrast, in the dynamic case, packets arrive over time; this arrival model is also referred to in the literature as asynchronous (see  (DBLP:conf/podc/Kowalski05, )). The literature adopts the terminology of arrival time or activation time to indicate when the corresponding participant begins executing a contention resolution protocol in order to send this packet. These times may be governed by a stochastic process (see the survey by Chlebus (chlebus2001randomized, )), or they may be scheduled by an adversary; in this survey, we focus on the latter setting.
Channel Characteristics. When a participant sends a packet, the multiple access channel provides feedback. The most basic feedback—assumed in all the literature we survey—occurs when, upon sending its packet, a participant learns whether or not its transmission was successful.
Another type of feedback concerns whether participants can detect collisions. Almost all of the literature on contention resolution assumes one of the following two types of feedback regarding collisions. The first is collision detection (CD): in any slot, each listening participant can differentiate between the case where the number of participants transmitting is 00 (i.e., an empty slot) or greater than $1$ (i.e., a collision). The second is no collision detection (no-CD): in any slot, a listening device cannot differentiate between the slot being empty versus containing a collision.
In many works, participants that listen to the channel also learn whether or not a slot contains a successful transmission. In such a setting with CD, the channel provides ternary feedback (sometimes referred to as “trinary feedback”) indicating whether the slot is empty, contains a successful transmission, or contains a collision. Similarly, in this setting, no-CD provides binary feedback. [^1] Other variants of CD appear in the wider literature involving multiple access channels (see  (chang:exponential, ; chang:exponential-jacm, )). Additionally, more exotic channel models have been considered; for example, De Marco and Kowalski (DBLP:conf/icdcs/MarcoK10, ) and Tsybakov (tsybakov:multiplicity, ) examine the case where channel feedback yields the number of senders.
Finally, some papers consider the case where more than one channel is available to participants; this is typically referred to as a multi-channel setting. Here, typically participants may select any one (and only one) channel to send on or listen to in any give slot. In many cases, a result involving a single channel may extend to the multi-channel setting without altering the core finding; however, there are cases where the use of multiple channels can admit significantly stronger results. Typically, this occurs when multiple channels are leveraged to achieve better latency or robustness to malicious interference, and throughout this survey, we highlight such results. In practice, multiple channels are available on several communication standards, such as IEEE 802.11 (WiFi), Bluetooth, and cellular standards (e.g., 4G and 5G LTE). Therefore, there is real-world motivation to push for results that make good use of more than one channel. However, in comparison to the single-channel setting, there are fewer results in the contention-resolution literature involving multiple channels, and so our survey is weighted more towards single-channel results.
Network Topology. A single-hop network is one where each participant can communicate with every other participant directly. In contrast, in a multi-hop network, a communication may need to pass between many participants until it reaches its intended recipient.
For the contention resolution problem, the type of topology may impact channel feedback. With CD in a single-hop network, all participants can detect the same channel feedback in any fixed slot. However, in a multi-hop network, it is common for different sets of participants to witness different channel feedback and, therefore, potentially update their respective states differently. Given this challenge, it is not surprising that the majority of work on contention resolution has focused the single-hop setting.
Global Knowledge. What information do all packets have about the system a priori? Having all packets agree on certain system parameters can significantly ease the challenge of contention resolution.
  • System-Size Knowledge. The exact number of the participants may be known a priori. Alternatively, an upper bound may be common knowledge. For example, in some problem formulations (see Section 3.2), only a subset of $k\leq n$ participants have a packet to send, where $n$ acts as an upper bound that is known or tightly-estimated a priori.
  • Unique Identifiers. Given $n$ participants, each may have a unique identifier (ID) taken from a set of $\Theta(n)$ possible IDs. This ID may be used by an algorithm to determine when the corresponding station attempts to send.
  • Global Clock. In the static case, all participants obtain their respective packet in the same slot $s$. This offers a common point for synchronization; that is, all participants may consider $s$ to have index 00. Thus, for the static case, there is the implicit assumption of a global clock. For dynamic arrivals, packets may be generated at participants at different times. In this case, if all participants agree on the index of the current slot, then they are said to have access to a global clock; otherwise, they are often said to have only a local clock.
1.1.3. Metrics.
In this section, we adopt the packet-centric terminology. In the static case, a natural performance measure is the number of slots until a set of packets succeeds; this number of slots is referred to as the makespan. We are typically interested in the set of all packets, but any deviations from this will be made clear from the context. For the static case, the ratio of the number of successful packets to the makespan is known as throughput, and this is another common notion of performance; note that throughput is a value between 00 and $1$.
For the dynamic case, we define a slot to be active, if at least one packet is present in the system during that slot. The above metrics remain valid when measured over active slots. However, at intermediate points in time, throughput may offer an overly-pessimistic measure of progress. For example, if a batch of $n$ packets arrive at the same time and start executing BEB, the throughput will be 00 for $\Omega(n/\log n)$ slots, despite the fact that the algorithm is progressing towards having many successful packet transmissions, which occurs after $\Theta(n)$ slots are executed.
To provide a more meaningful metric, Bender et al. (bender2020contention, ) propose the notion of “implicit throughput”. The implicit throughput at time $t$ is defined as $N_{t}/A_{t}$, where $N_{t}$ is the total number of packets that arrive at or before time $t$, and $A_{t}$ is the total number of active slots at or before time $t$. Returning to the example with a batch of $n$ packets, we note that the implicit throughput is large—specifically, $n$—in the first slot, and is $\Omega(\log n)$ over the first $\Omega(n/\log n)$ slots. The implicit throughput will continue to decrease, reaching $\Theta(1)$. Then, ultimately, the implicit throughput will be $\Theta(1/\log n)$ once all packets have succeeded, which matches the standard measure of throughput for BEB, as mentioned in Section 1. An algorithm that guarantees an implicit throughout of $\Omega(1)$ over all slots performs well, since the number of active slots so far is never asymptotically larger than the number of packets that have arrived. Furthermore, as exemplified above, at points in time where the standard notion of throughput is meaningful—such as when there are no packets in the system—the two metrics are equal.
Continuing with the dynamic case, another metric is the number of slots it takes for a participant to successfully transmit its packet starting from its activation time; this is latency. We are interested in the maximum latency over all participants. Here, we can extend throughput to be the number of packets over the maximum latency.
In much of the early literature on the dynamic case, packets are generated according to a stochastic process, and efficiency is tied to the packet arrival rate; that is, the number of packets that are newly generated (arrive) in each slot and held by (some subset) of participants. For any participant, its unsent packets are stored in a queue for eventual transmission on the multiple access channel, and this number of unsent packets is referred to as the queue size. Given a packet arrival rate, the goal is to guarantee that every queue has bounded size. The higher the rate that can be tolerated, the more efficiently the algorithm allows the network to digest new packets. In Section 3.1, we expand on this notion, which is generally referred to as stability.
Given the popularity of small, battery-powered wireless devices in modern networks, many contemporary results focus on energy efficiency. The energy consumed by a participant is typically measured as the number of slots that the participant accesses the channel, as sending and listening often dominate the operational costs of such devices (SINHA201714, ; feeney2001investigating, ; polastre:telos, ; wang2006realistic, ). We refer to this as the number of channel accesses.
Robustness to malicious interference—often referred to as jamming—is becoming increasingly important, with several well-publicized attacks in recent years (nicholl, ; FCC:humphreys, ; marriott, ; jamming-church, ). Contention-resolution algorithms that rely on channel feedback are especially vulnerable to adversarial noise. We refer to this aspect as attack resistance and, for each relevant result, we describe the type of attack that can be tolerated.
1.1.4. Algorithm Design.
There is some common terminology used in the literature to characterize different classes of contention resolution algorithms. An obvious division is with respect to deterministic versus randomized algorithms, where the former is characterized by the absence of random bits being used in the execution. Deterministic solutions are appealing from a practical standpoint in that their guarantees hold even when the number of participants and/or packets is small, admitting no error probability. In trade, they tend to provide sub-optimal asymptotic results in many settings, compared to randomized solutions, as we shall see.
An adaptive algorithm knows the full history of channel feedback and can use this information to inform future actions. Messages other than the sending of the packet itself are allowed; for example, such messages might pertain to estimates of the system size or coordination information. Sometimes this content is referred to as control bits (hradovich2020contention, ), and they are often used to facilitate coordination (see (chlebus:adversarial, )). In contrast, the only feedback that might alter the execution of a non-adaptive algorithm is a participant learning that its packet has been successfully transmitted, after which the participant terminates (doi:10.1137/140982763, ). There is a less-common notion of strong non-adaptiveness, where the execution of a participant is not even altered by the outcome of a transmission (again, see (doi:10.1137/140982763, )).
Figure 1. An illustration of features used to structure our survey of the literature. Larger boxes correspond to general features, while smaller boxes indicate those that are more fine-grained.
Finer distinctions are sometimes used. A full-sensing algorithm has the same characteristics as an adaptive algorithm, but no control bits are allowed (chlebus2006adversarial, ; hradovich2020contention, ). Finally, an algorithm where packets can use acknowledgments (implicitly received after succeeding) to terminate is said to be acknowledgement-based (DBLP:conf/icdcs/MarcoKS19, ); otherwise, if packets continue executing after their respective successes, the algorithm is non-acknowledgement-based. [^2] This terminology varies slightly in the literature. In the stochastic dynamic model, an acknowledgement-based algorithm depends on the packet’s own transmission history (Goldberg-notes-2000, ). In adversarial queueing theory (AQT) (see Section 3.1), it means a packet is sent according to a schedule based on its ID and the current round (chlebus2006adversarial, ; DBLP:conf/icdcs/MarcoKS19, ; hradovich2020contention, ). The latter definition incorporates a dependency on IDs, since much of the work in the acknowledgement-based setting involves deterministic algorithms, which require unique IDs (see Section 3.2 for more discussion).
These classifications can be useful, since they indicate the strength of the underlying assumptions being made. For example, an algorithm that is randomized and full-sensing makes stronger assumptions than an algorithm that is deterministic and acknowledgement-based. A device running the former algorithm must have a pseudorandom number generator and be able to receive feedback (beyond learning of its own success) on the channel. In contrast, the latter algorithm can function even when channel sensing is unavailable—either due to the nature of the channel or a participant’s hardware constraints—and the resulting protocol may be less complex in its implementation by eschewing randomness and knowledge of transmission history.
However, we note that this classification can be complicated to apply in certain cases; for example, in a multi-channel setting, a device may not have the full history of feedback on all channels. Therefore, in situations where the classification is clear, we present this information; otherwise, we omit it.
1.2. Survey Organization
We use the attributes discussed above and illustrated in Figure 1 to structure our review of the literature. In Section 2, we begin with contention-resolution algorithms for the static case. This provides an introduction to the important metric of makespan, as well as the use of randomization in designing contention-resolution algorithms. Next, in Section 3, we move to dynamic model of packet arrivals, focusing on the case when arrival times are dictated by an adversary. Finally, in Section 4, we address additional adversarial behavior in the form of malicious interference on the channel.
Each section has the following associated components:
  1. (1)
    A specific result that we feel provides some intuition for an important aspect of the results surveyed. This helps us to highlight (i) those challenges posed by the contention-resolution problem being addressed, and (ii) one way in which these challenges can be overcome, which in turn, allows us to contrast against some other surveyed solutions.
  2. (2)
    A discussion portion, where we highlight the main themes of the literature surveyed, along with some potential directions for future work.
  3. (3)
    A table that offers a synopsis of the results in each section by organizing the results along common, pertinent attributes from Section 1.1. A brief summary of the result is also provided, which can be useful for reference purposes. Due to space constraints, these tables can be found in our electronic supplement.
2. The Static Case
We begin our survey with the static case for packet arrivals. This motivates a well-known class of randomized, non-adaptive backoff algorithms, with origins in decades-old networking protocols; we explore this topic in Section 2.1. Additionally, the static case also allows us to introduce deterministic algorithms, which are newer additions to the literature; we review these results in Section 2.2. Finally, we discuss the closely-related wakeup problem in Section 2.3.
2.1. Backoff Algorithms
Arguably, the most well-known backoff algorithm is binary exponential backoff (BEB), which was introduced by Metcalf and Boggs (MetcalfeBo76, ), and briefly discussed in Section 1. BEB proceeds in disjoint contention windows (CWs) consisting of consecutive slots. In each CW, a participant with a packet selects a slot uniformly at random to attempt a transmission. If successful, another packet may be processed in the same fashion, if one exists. Else, transmission is attempted in the next CW, which is double the size of the current CW. This is illustrated in Figure 2, along with pseudocode.
While originally intended for use with Ethernet over bus networks (stuck1984introduction, ), BEB has since found application in a range of domains such as wireless communication (zhou2019singletons, ; sun:backoff, ; 6859627, ; saher:log, ), transactional memory (herelihy:transactional, ; scherer:advanced, ), concurrent memory access (Ben-DavidB17, ; 80120, ; mellor-crummey:algorithms, ), and congestion control (mondal:removing, ). [^3] A ubiquitous example is WiFi, where many of the standards use BEB as a core mechanism for governing how devices gain access to a wireless channel. In these standards, the initial window size is typically some constant larger than $1$, and the maximum window size is limited to a value just over $1000$. Despite this widespread use, the makespan for BEB is sub-optimal (see Bender et al. (bender:adversarial, )); specifically, for a batch of $n$ packets that begin executing BEB at the same time, the makespan is $\Theta(n\log n)$. To understand why, we need to become familiar with the notion of contention.
Figure 2. Pseudocode for BEB and an illustration of the first few contention windows. Time moves from left to right.
Developing Intuition.
In order to better understand why BEB performs poorly, we consider a class of Bernoulli-backoff algorithms, where each packet sends with probability $1/w$ in each slot of a window of size $w$. Note that this differs subtly from the approaches discussed in Section 2.1, where a single slot is chosen uniformly at random in which to send. For Bernoulli-backoff algorithms it is easy to analyze the useful notion of contention (richa:jamming4, ; ChangJP19, ; BenderFGY19, ; DBLP:conf/spaa/AgrawalBFGY20, ; bender:how, ). Let $p_{i}(t)$ denote the probability that packet $i$ attempts to broadcast in slot $t$. We define the contention in slot $t$ to be $C(t)$ $=\sum_{i}p_{i}(t)$; that is, the sum of the sending probabilities for all packets in the system at time $t$.
The notion of contention characterizes the probability that some packet is successful in slot $t$; denote this quantity by $p_{\mbox{\tiny suc}}(t)$. In particular, if $p_{i}(t)\leq 1/2$, which is true if we set the initial window to be size at least $2$, then the following relationship holds:
(1) $\frac{C(t)}{e^{2C(t)}}$ $\leq p_{\mbox{\tiny suc}}(t)\leq\frac{2C(t)}{e^{C(t)}}.$
Like the temperature of the porridge in the Goldilocks fable, the contention can be “too high” or “too low” for packets to succeed. For example, consider the Bernoulli equivalent of BEB. Initially, the contention is high. In particular, $C(t)=\Omega(\ln n)$ which implies that $p_{\mbox{\tiny suc}}(t)$ is polynomially small in $n$ until a window of size $\Theta(n/\ln n)$. We highlight that, while we use Equation 1 to gain intuition about the static case, these bounds on contention will also be useful later on when discussing the dynamic case (Section 3).
When the window size reaches $\Theta(n)$, $C(t)=\Theta(1)$ and so $p_{\mbox{\tiny suc}}(t)=\Theta(1)$. At this point, the contention is “just right” and many packets succeed. Unfortunately, since the window size grows aggressively under BEB, the contention rapidly decreases, and so, any remaining packets require many slots before they succeed. This behavior is what leads to suboptimal makespan, and it begs the question: is there a backoff algorithm that yields the asymptotically optimal makespan of $O(n)$ or, equivalently, $\Omega(1)$ throughput? As a first step towards answering this question, we consider backoff algorithms where the growth of contention windows differs from BEB.
A Generic Backoff Algorithm $i$: index of current window for $i\geq 0$. $w_{i}$: size of window $i$. $f(i)$: window-scaling function. In current window $i$, a station with a packet to transmit does the following until successful: Attempt to transmit in a slot chosen uniformly at random from $w_{i}$. If the transmission failed, then wait until the end of the window and $w_{i+1}\leftarrow(1+f(i))w_{i}$.
Figure 3. Generic backoff with monotically-increasing windows. For LB, LLB, BEB, and PB, the initial window size is a constant, and $f(i)=1/\lg w_{i},1/\lg\lg w_{i}$, $1$, and $(1+1/i)^{c}-1$, respectively. For FB, $f(i)=0$ and the initial window size is $\Theta(n)$.
General Monotonic Backoff
The makespan of a backoff algorithm depends on the amount by which successive CWs increase in size; for BEB, this is simply a factor of $2$. Several results examine different values (Ni:survey, ; 1543687, ; sun:backoff, ; 6859627, ; saher:log, ), presenting a mixture of empirical and analytical results when the network is saturated; that is, when each participant always has a packet to send (bianchi1998ieee, ). In contrast, the static model of arrivals captures a type of bursty traffic, which turns out to be more challenging. In this setting, Bender et al. (bender:adversarial, ) derive makespan results for BEB and several closely-related variants: Log-Backoff (LB), LogLog-Backoff (LLB), and Fixed Backoff (FB). These variants operate similarly to BEB: when the current CW of size $w$ is finished, the next CW has size $(1+f(w))w$, where $f(w)=1/\lg(w),1/\lg\lg(w),$ $1$, and 00, for LB, LLB, BEB, and FB, respectively. [^4] We use $\lg(x)$ to denote the logarithm base 2 of $x$. In each CW, a participant randomly selects a single slot in which to send, and other than learning of success or failure, no channel feedback is used by the algorithms; all of these algorithms are acknowledgement-based.
The pseudocode for these backoff algorithms is presented in Figure 3. To the best of our knowledge, the initial CW sizes for LB, LLB, and FB are not specified exactly in the literature; $\Theta(1)$ will suffice for $LB$ and $LLB$. A conspicuous requirement of FB is that $n$ must be known to within a constant multiplicative factor in order to set the CW size. A number of prior works show how to get an estimate of $n$ (DBLP:journals/jacm/GreenbergFL87, ; bender:contention, ; cali:dynamic, ; cali:design, ; bianchi:kalman, ).
In terms of performance, for a batch of $n$ packets, the makespan for BEB, LB, FB, and LLB is $\Theta(n\log n)$, $\Theta\left(\frac{n\log n}{\log\log n}\right)$, $\Theta\left(n\log\log n\right)$, and $\Theta\left(\frac{n\log\log n}{\log\log\log n}\right)$, respectively (bender:adversarial, ; zhou2019singletons, ). Bender et al. also analyze polynomial backoff (PB), where the $i^{\mbox{\tiny th}}$ window has size $\Theta(i^{c})$ for some constant $c>1$. The window scaling function is slightly different here; ignoring constants, we solve for $f(i)$ in $(i+1)^{c}=(1+f(i))i^{c}$, implies that $f(i)=((1+i)^{c}/i^{c})-1=((i+1)/i)^{c}-1$. For PB, Bender et al. show that the makespan is $\Theta(n/\log n)^{1+(1/c)}$.
Interestingly, all of these algorithms suffer from a throughput that tends to zero. This observation is striking for a couple reasons. First, it illustrates the challenges raised by the static case. In particular, the throughput for PB when $c=2$, known as quadratic backoff, is shown by Sun and Dai (sun:backoff, ) to be competitive with BEB under a saturated network. In contrast, under static arrivals, quadratic backoff has asymptotically inferior makespan.
Second, from a theoretical perspective, does a backoff algorithm exist that gives optimal throughput under the static case? Bender et al. (bender:adversarial, ) show that, over the class of monotonic windowed backoff algorithms, LLB achieves the asymptotically-best makespan  (bender:adversarial, ). In other words, a non-monotonic approach is required to achieve $\Theta(n)$ makespan or, equivalently, $\Theta(1)$ throughput. This question of asymptotic optimality is what we pursue next.
Understanding Performance.
To develop intuition for monotonic backoff algorithms, we start by asking: why does (the Bernoulli version of) LB fare better than BEB? Again, initially the high contention prevents packets from succeeding. Indeed, the high contention persists for longer under LB given that the window size grows far more slowly compared to BEB. However, once (near) constant contention is achieved, packets succeed. Critically, packets stay within the Goldilocks zone for longer since, again, the window size increases more slowly relative to BEB. Finally, contention drops off more gradually, allowing any straggling packets to succeed within fewer slots than under BEB.
The same qualitative behavior is true of LLB, which grows its window size faster than LB, but more slowly than BEB. This allows LLB to strike the asymptotically optimal balance between the number of wasted slots spent in either high-or-low-contention windows, and those slots with constant contention.
Why doesn’t FB perform better? After all, it avoids the high-contention slots by skipping directly to a window of size $\Theta(n)$ where the probability of success is constant. Here, again, contention tells the story. While initially many packets succeed under FB, the contention quickly decreases, and this inflicts an $O(\log\log n)$-factor over the optimal makespan.
Ultimately, for the static case, the above algorithms fall short due to their inability to provide constant contention over $cn$ slots for some sufficiently large constant $c$. Next, we examine algorithms that do satisfy this property.
Achieving Asymptotically Optimal Makespan
Based on our observations in Section 2, any algorithm that hopes to achieve $O(n)$ makespan must incorporate a mechanism for providing constant contention over $\Omega(n)$ slots. Sawtooth-Backoff (STB) (GreenbergL85, ; Gereb-GrausT92, ) accomplishes this; it is non-monotonic algorithm that executes over a doubly-nested loop. The outer loop sets the current window size $w$ to be double the one used in the preceding outer loop; this is like BEB. Additionally, for each such window, the inner loop executes over $\lg w$ windows of decreasing size: $w,w/2,w/4,...,1$. For each such window, a slot is chosen uniformly at random for the packet to transmit; this is the “backon” component of STB. When the outer loop achieves a window of size $cn$ for a sufficiently large constant $c>0$, with high probability all packets will succeed in the corresponding inner loop. Intuitively, this is because the backon component maintains constant contention in each slot of the inner-loop windows (until all packets succeed). For a single batch of $n$ packets, the makespan of STB is $O(n)$ (Gereb-GrausT92, ; GreenbergL85, ; bender:adversarial, ; zhou2019singletons, ).
A variant called Truncated Sawtooth-Backoff (TSTB) (bender:contention, ) executes identically to STB except that, for some constant $c>0$, the inner loop executes over $\lg(c\lg w)$ windows of decreasing size: $w,w/2,w/4,...,\max\{\lfloor w/c\lg n\rfloor,1\}$. Another asymptotically-optimal result is given by Anta et al. (FernandezAntaMoMu13, ; fernandez-anta:unbounded, ) with their algorithm One-fail Adaptive, which succeeds with probability $1-O(1/n)$, again for the case with no-CD and no knowledge of system size. The authors tightly analyze both One-fail Adaptive and STB, yielding respective makespan guarantees of $2(\delta+1)n+O(\log^{2}n)$ and $4(1+1/\delta)n$, where $\delta$ is a small positive constant.
Anta and Mosteiro (Anta2010, ) look at a formulation of the static case—referred to as static $k$-selection (see Section 3.2)—where an unknown $k$ out of $n$ participants must each succeed. For an error parameter $\epsilon\leq 1/(n+1)$, the authors give a randomized algorithm that with probability $1-O(\epsilon)$ has makespan $O(k+\log^{2}(1/\epsilon))$; for consistency, we report the case for $k=n$ in Table 1 of our electronic supplement.
For the special case where $n=2$, Wang (dingyu:optimal, ) provides exact, optimal results. Specifically, the average latency until both participants succeed, the latency until the first participant succeeds, and the latency until the second (last) device succeeds are shown to have expected values, respectively, of $\sqrt{3/2}+3/2$, $2$ and approximately $3.33641$.
These results settle the question of optimal makespan for the static case. However, an obvious question is whether this optimality holds when packets may arrive over time, and we explore this issue in Section 3.
2.2. Deterministic and Hybrid Approaches
In contrast to randomized backoff, there are a handful of results that propose deterministic approaches for what is, essentially, the case of static-arrivals. Specifically, packets may arrive over time. However, once a collision occurs, those participants whose packets collided must execute a “conflict resolution” algorithm, while all other participants learn about the collision via channel feedback and refrain from transmitting; that is, no other packets are sent until this conflict is resolved.
For these results, there are $k\geq 2$ participants, each with a unique ID in the range $\{1,...,n\}$. Komlos and Greenberg (komlos:asymptotically, ) show the existence of a non-adaptive, deterministic algorithm that, with knowledge of $k$ and $n$, has a makespan of $O(k+k\log(n/k))$; their approach requires that each participant $u$ only adapts to channel feedback during slots in which $u$ transmits. Related to this, in the static case, Kowalski (DBLP:conf/podc/Kowalski05, ) proves the existence of a non-adaptive deterministic algorithm with $O(k\log(n/k))$ makespan, and also gives an explicit construction for a non-adaptive algorithm with $O(k\,\texttt{polylog}(n))$ makespan. These results stand in contrast to prior work by Capetanakis (capetanakis:generalized, ) and Tsybakov and Mikhailov (tsybakov1978free-russian, ; tsybakov1978free, ) who give deterministic algorithms with the same asymptotic makespan without needing to know $k$; however, their algorithms require that stations receive feedback even in slots where they themselves are not transmitting. Related to these results, Greenberg, Flajolet, and Ladner (DBLP:journals/jacm/GreenbergFL87, ; DBLP:conf/focs/GreenbergL83, ) give a randomized algorithm to estimate the number of stations that need to send a packet; this relies on collision detection. This is combined with prior deterministic approaches (capetanakis:generalized, ; hayes:adaptive, ; tsybakov1978free-russian, ; tsybakov1978free, ; fayolle1982capacity, ; hofri1984stack, ), yielding several hybrid algorithms, and the best expected makespan is at most $2.1338n+o(n)$.
2.3. The Wakeup Problem.
We introduce the wakeup problem, which is described as follows. There are $n$ participants, of which some are active (or “awake”), and the remainder are inactive (or“sleeping”) with only the ability to listen. Each awake participant has a packet to send, and upon the first success, all sleeping participants “wake up” and the problem is solved. Typically, the metric of interest is the number of rounds until the first success. Thus, contention resolution differs from wakeup: it requires all participants to succeed, while wakeup hinges on just the first success.
Despite this difference, historically the wakeup problem has considerable overlap in the literature with contention resolution, and is sometimes referred to by the same name (e.g., (fineman:contention, ; fineman:contention2, ; gilbert2021contention, ; DEMARCO20171, ; de2013contention, )). To continue the naming complications, in the static case where the set of active participants are fixed from the start, the wakeup problem is sometimes referred to as leader election (LE).
Static wakeup and LE may be viewed as a natural starting point for solving contention resolution, as repeatedly applying them over all $n$ packets offers a solution, albeit an inefficient one. A more meaningful connection can be drawn between static wakeup and contention resolution: wake-up protocols often produce an estimate for the $n$, the size of the system, which may be used to bootstrap a contention resolution protocol. More specifically, many algorithms are “uniform” or ‘fair”: in any fixed round, each participant sends packets with the same probability (independently of the prior communication history). For such algorithms, the reciprocal of the sending probability when the first success occurs may yield an accurate estimate of $n$; obtaining system-size knowledge is a well-known component of contention resolution (e.g., (jurdzinski:energy, ; chang:exponential-jacm, ; gilbert2021contention, )). For instance, such an estimate can be used to set the first window size in algorithms such as FB and STB/TSTB (recall Section 2.1), bypassing much of the backing off that would otherwise be required, and speeding up the runtime by a constant factor.
For static wakeup, makespan refers to number of slots until the first success (recall Section 1.1.3). An early result on the wakeup problem is by Willard (willard:loglog, ), who considers the case with CD. Here, Willard establishes a lower bound on uniform algorithms—for any fixed round, each participant sends packets with the same probability (independently of the prior communication history)—showing that the expected makespan is at least $\lfloor\lg\lg n\rfloor-O(1)$, and an upper bound of $\lg\lg n+O(1)$ for expected makespan is also given.
In the case of no-CD and where $n$ is known, Kushilevitz and Mansour (kushilevitz1998omega, ; kushilevitz1993omega, ) prove a lower bound of $\Omega(\log n)$ expected makespan; this applies even to unfair algorithms. An asymptotically-matching upper bound on expected makespan, and a bound of $O(\log^{2}n)$ w.h.p. is achieved by Bar-Yehuda et al. (bar1992time, ). In subsequent work that unifies many prior lower-bound arguments, Newport (newport2014radio, ; newport:radio-journal, ) also shows $\Omega(\log n)$ expected makespan and $\Omega(\log^{2}n)$ makespan w.h.p. for any randomized algorithms; these bounds hold even with multiple channels and/or CD. For algorithms that succeed with probability at least $1-\epsilon$, this lower bound was strengthened to $\Omega(\log n\,\log(1/\epsilon))$ by Farach-Colton et al. (farach-colton:lower-bounds, ). We note a couple things. First, these lower bounds are dramatically different from the lower bound by Willard (willard:loglog, ), where CD is available. Second, Jurdziński and Stachowiak (jurdzinski2002probabilistic, ) provide an algorithm with $O(\log n\,\log(1/\epsilon))$ makespan so long as at least one of the following is true: devices have unique labels, or $n$ is known.
In a setting with $\mathcal{C}>1$ channels and CD, Fineman et al. (fineman:contention2, ) give an algorithm guaranteeing w.h.p. $O((\log n/\log\mathcal{C})$ $+(\log\log n)(\log\log\log n))$ makespan for the case where any number of participants may start in the active state, along with an algorithm guaranteeing w.h.p. $O((\log n/\log\mathcal{C})+(\log\log n))$ makespan in the restricted case where at most two participants become active; this second case matches the lower bound provided by Newport (newport2014radio, ; newport:radio-journal, ). The authors also note that, while their result is described for the static case, it holds for the dynamic case.
The signal-to-interference-plus-noise ratio (SINR) model has been considered in the context of several fundamental wireless-communication problems (for details, see (moscibroda2006complexity, ; jurdzinski2012distributed, ; jurdzinski2013distributed, ; moses2020deterministic, ; jurdzinski2015cost, ; JurdzinskiR17, )). Informally, under the SINR model, the ratio of (i) the power of an incoming transmission to (ii) the sum interference from other devices plus the floor noise, must exceed a threshold value in order for the transmission to succeed. In this setting, Fineman et al. (fineman:contention, ) shows how to solve wakeup with $O(\log n+\log R)$ makespan w.h.p., where $R$ is ratio between the longest and shortest link. The authors point out that, in many practical settings, $R$ is at most a polynomial in $n$, which yields an $O(\log n)$ result. [^5] This ratio appears in other related results under the SINR model, such as interference scheduling or link scheduling (e.g., (fanghanel2011improved, )). The authors also show a general lower bound (e.g., it is not restricted to the class of fair algorithms) of $\Omega(\log n)$.
Finally, we note a recent result by Gilbert et al. (gilbert2021contention, ), that examines the implications of system-size “predictions”; that is, the algorithm is given access to the probability distribution defined over the possible values of $n$. The authors give lower and upper bounds for both deterministic and randomized solutions, for the CD and no-CD cases. Interestingly, in addition to new bounds, this work recovers some lower bounds, including the $\Omega(\log\log n)$ lower bound by Willard (willard:loglog, ) mentioned above.
Summary. For static contention resolution, many randomized algorithms have been proposed, with STB providing one of the strongest with-high-probability results: asymptotically-optimal makespan with no-CD, no knowledge of $n$, and only $O(\log n)$ channel-access attempts.
In contrast, purely deterministic approaches cannot achieve the asymptotically-optimal makespan without some additional assumptions regarding CD, knowledge of $n$, or a unique labeling of participants. We see a similar discrepancy when we consider adversarial dynamic arrivals, which we explore next.
In the context of network traffic, we highlight the underlying assumption that each transmission of a packet is accommodated by a single slot. Unsurprisingly, BEB also has sub-optimal makespan for the more general case with heterogeneous packet sizes, where a packet of size $s$ must have sole access to the channel for $s$ consecutive slots in order to succeed (bender:heterogeneous, ). Along similar lines, in the wireless domain, it has been argued that packet size should be incorporated into the definition of makespan, since in WiFi networks collisions may consume time proportional to packet size (anderton:is, ; anderton:windowed, ).
For static wakeup, under each of CD or no-CD, asymptotically matching upper and lower bounds are known. It is interesting to compare these results to those of static contention resolution. In particular, while the makespan for contention resolution is higher, the optimal average per-packet time to success is $\Theta(1)$, which is lower than even the CD case for wakeup. As we will see, this is also a common theme in the dynamic case with randomized approaches.
3. Dynamic Arrivals
What can be said about the case where packets do not arrive and succeed in disjoint batches, but rather are injected into the system over time? A large body of literature addresses contention resolution where arrival times are governed by a stochastic process, typically Bernoulli or Poisson. Informally, with a stochastic setting, the aim is to show that packets do not grow without bound at each sender, but are successfully delivered at a pace sufficient to keep up with the rate that new packets are generated. For the interested reader, we highlight the survey by Chlebus (chlebus2001randomized, ), which addresses in depth the literature in this area.
However, modeling of arrival times via a stochastic process may not always be justified. In particular, bursty traffic is common in practice (williamson:internet, ; wilson1991high, ; paxson:wide-area, ; Teymori2005, ; yu:study, ), and its impact has been examined (Ghani:2010, ; sarkar:effect, ; canberk:self, ; bhandari:performance, ). Indeed, the static case discussed previously in Section 2 illustrates the impact on makespan by disjoint bursts of packets. This has led the research community to consider more general arrival models.
To this end, adversarial queuing theory (AQT) (borodin2001adversarial, ; cruz-one, ; cruz-two, ; andrews2001universal, ) has been used to model packet arrivals times, and many contention-resolution results fit within the AQT framework; that is, they address adversarial models that are parameterized by rates of packet injection and burstiness. Here, a key metric is stability, which we discuss in Section 3.1.
Another large body of the contention-resolution literature eschews the parameterization of AQT and instead considers a completely arbitrary schedule for when packets arrive. Here, we treat separately the literature on deterministic and randomized approaches, in Sections 3.2 and 3.3, respectively.
3.1. AQT and Contention Resolution
In this setting, each participant is sometimes associated with a queue into which packets may be injected by an adversary at arbitrary times. Other results are presented in a queue-free model, where each packet has a corresponding station that “dies” (i.e., becomes irrelevant) after the packet is successfully transmitted (see discussion by Goldberg (Goldberg-notes-2000, ; anantharamu2010deterministic, )). In either case, packet arrival times are parameterized by an injection rate $\rho$, where $0<\rho\leq 1$, and a non-negative integer, $b$, that denotes burstiness.
Several adversarial models appear in the AQT literature; see Borodin et al (borodin2001adversarial, ) for an in-depth discussion. In the context of contention resolution, there are two prominent models. The leaky-bucket adversary is defined such that, for any $t$ consecutive slots, at most $\rho t+b$ packets may be placed in the queue of any participant (cruz-one, ). For a contiguous segment of rounds, and any positive integer $w$, a window adversary of type $(\rho,w)$ can inject $\rho w$ packets in each contiguous segment of $w$ rounds into any set of stations; $w$ is referred to as a window.
Performance is typically measured by the growth of queues at each participant. As defined by Andrews et al. (andrews2001universal, ), an algorithm is stable on a network despite an adversary if, starting from a state where there are no packets, the total number of packets is always bounded by a constant. An algorithm is universally stable if it is stable against every adversary with $\rho<1$, for any network (andrews2001universal, ). Finally, in the contention-resolution literature, the notion of strongly stable arises in the context of a window adversary: an algorithm is strongly stable if, at any time, the total number of queued packets is $O(\rho w)$ (chlebus2006adversarial, ; chlebus:adversarial, ).
Another important metric is (packet) latency, which is often defined as the maximum time that any packet stays in the queue prior to being successfully transmitted. We note that, while this is described in the context of queues, this aligns with the metric as described in Section 1.1.3.
Figure 4. Depiction of how exponential backoff is unstable for a packet arrival rate of $\Omega(\ln\ln n/\ln n)$.
Developing Intuition.
Given the prominence of backoff algorithms, one may wonder about their performance in the AQT setting. Given the suboptimal makespan of popular algorithms like BEB in the static case, we should be skeptical of their performance when packets arrive dynamically at a high rate. To further this intuition, we sketch an argument that shows the instability of an algorithm BEB’, which is similar to BEB. Specifically, for each slot $i\geq 1$ starting with the slot that packet $u$ injected into the system, $u$ sends with probability $1/i$. Note that this is similar to the Bernoulli-backoff algorithm (recall Section 2.1), where the sending probability per slot proceeds as $1/2,1/2,1/4,1/4,1/4,1/4,1/8,1/8,...$ etc. By comparison, BEB’ proceeds as $1,1/2,1/3,1/4,1/5,...$, and so at any given slot since a packet is injected, the sending probabilities under the Bernoulli-backoff algorithm and BEB’ differ by only a constant. [^6] We highlight that this is very similar to the algorithm Decrease Slowly given by Jurdziński and Stachowiak (jurdzinski2005probabilistic, ; jurdzinski2002probabilistic, ); see the proof of Theorem 8 in (jurdzinski2005probabilistic, ) for a formal analysis.
We can now describe why BEB’ is unstable for an arrival rate of $\Omega(\ln\ln n/\ln n)$; Figure 4 complements our discussion. Consider a window $W$ of $T=\Theta(n\ln n/\ln\ln n)$ slots. In the first slot, a single burst of $n/2$ packets arrives. This drives up the contention to $\Omega(n)$ in this slot. None of these packets is likely to succeed over the next $n/\ln n$ slots; to see this, note that in the $i$-th slot measured from this first slot is sending with probability $1/i$, yielding a contention of $\Omega(\ln n)$ in each of these $n/\ln n$ slots, and so the probability of success is $O(1/{\texttt{poly}({\ln n})})$ by Equation 1.
In addition to the burst, there is a steady stream of $n/2$ packets: a single packet is injected every $\Theta(\ln n/\ln\ln n)$ slots measured after the first slot. Given our discussion in Section 2.1, by itself, this stream poses no problem for BEB’. Specifically, by the lower bound in Equation 1, we expect each packet to succeed within a small constant number of slots prior to the next packet arriving and, therefore, contention remains low.
However, given the burst of packets, the packets in the stream face $n/\ln n$ slots where contention is $\Omega(\ln n)$. Consequently, rather than succeeding, these packets accumulate in the system and, over these $n/\ln n$ slots, their contribution to contention is $\Omega(\ln\ln n/\ln n)\sum_{i=1}^{n/\ln n}1/i=\Omega(\ln\ln n)$ by the sum of a harmonic series. By Equation 1, the probability of success is $O(1/{\texttt{poly}({\ln n})})$, and thus the expected number of slots for a packet to succeed is $\Omega({\texttt{poly}({\ln n})})$.
To summarize the above situation, we expect that for all $n$ packets to succeed $\Omega(n\cdot{\texttt{poly}({\ln n})})$ slots are required, which exceeds $T$. Consequently, for every such interval $T$, more and more packets accumulate in system over time. Note that this instability occurs for a rate that is $n/T=\Omega(\ln\ln n/\ln n)$. The above argument originates with Bender et al. (bender:adversarial, ), who use it to establish bounds on the arrival rate for which several backoff algorithms are unstable.
Overview of Literature. As mentioned above, the result by Bender et al. (bender:adversarial, ) explores stability of backoff algorithms under windowed adversaries in the queue-free model. BEB is shown to be stable for $\rho=d/\log n$ for some constant $d>0$, and unstable for $\rho=\Omega(\log\log n/\log n)$ They also show that LLB is stable for $\rho=d^{\prime}/((\log n)\log\log n)$ for some constant $d^{\prime}>0$, and unstable for $\rho=\Omega(1/\log n)$. Overall, traditional backoff algorithms appear to perform sub-optimally, although this is not surprising given the corresponding makespan results for the static case.
The remaining approaches in the literature are deterministic. Chlebus et al. (chlebus2009maximum, ; chlebus2007stability, ) consider adversaries where $\rho=1$; that is, one packet is injected per slot. The main positive result is a deterministic, adaptive protocol that guarantees against leaky-bucket adversaries that $O(n^{2}+b)$ packets are queued in the system at any given time, where $n$ is the number of participants. Here, $n$ is known to all participants and each ID is unique in the range $[1,...,n]$; this allows participants to schedule via passing a token, such that no collisions occur (notably, the positive results in this paper do not require collision detection). This result is shown to be asymptotically optimal in the sense that a leaky-bucket adversary with burstiness $2$ can force $\Omega(n^{2})$ total packets to be queued.
Several impossibility results are also given—holding regardless of whether or not collision detection is available—that illustrate how even minor burstiness can prevent stability. Specifically, no full-sensing protocol can be stable (i) for at least two participants against a leaky-bucket adversary with burstiness $2$, (ii) or for at least three participants against a window adversary of burstiness $2$; and (iii) no acknowledgment-based algorithm is stable in a system of at least two participants against the window adversary of burstiness $2$.
Additional impossibility results are made in the context of three natural properties. First, an algorithm is fair if each packet is transmitted successfully within $O(b/\rho)$ slots from the time it is injected. [^7] Note that this is a different notion of fairness than described in Section 2.3. The authors show that no algorithm can be fair and stable for any number of at least two participants against a leaky-bucket adversary of burstiness $2$. Second, an algorithm is withholding the channel if, upon successfully transmitting a packet, the station keeps sending until all of its pending packets are exhausted. No algorithm that withholds the channel can be stable for three stations against a leaky-bucket adversary of burstiness $2$. Third, an algorithm is said to be retaining if after a participant successfully transmits a packet, and if any packets arriving subsequent to this event are injected only at other stations, the participant eventually empties its queue. No retaining protocol is stable in a system of at least four stations against a window adversary of burstiness $2$. Chlebus et al. show that this implies no protocol that is fair can also be stable in a system of at least four stations against the window adversary of burstiness $2$.
In contrast to the above, Chlebus et al. (chlebus2006adversarial, ; chlebus:adversarial, ) study deterministic algorithms under the same model as Chlebus et al. (chlebus2009maximum, ; chlebus2007stability, ), but where $\rho<1$. They show that strong stability is impossible for a rate $\omega(1/\log n)$, with or without collision detection. When collision detection is available, the authors give a full-sensing algorithm that is fair and stable for $\rho=O(1/\log n)$, and show that fairness is not possible for rates that are $\omega(1/\log n)$. When collision detection is not available, the authors give an existence proof for a stable and fair algorithm under a much smaller rate of $O(1/n\log^{2}n)$, and they give an explicit algorithm that achieves these properties for a rate of $O(1/n^{2}\log n)$. Finally, it is shown that no acknowledgement-based algorithm can be stable for a rate exceeding $3/(1+\lg n)$.
Anantharamu et al. (anantharamu2010deterministic, ) and Aldawsari et al. (aldawsari:adversarial, ) design and analyze several deterministic algorithms that employ either a round-robin or token-passing approach for obtaining access to the channel. The authors derive bounds on latency under the leaky-bucket adversary model as a function of $\rho$, for $\rho<1/(2\log n)$ and general $b$. Note that this restriction on $\rho$ should not surprise us, given the negative results on strong stability by Chlebus et al. (chlebus2006adversarial, ; chlebus:adversarial, ) and prior results on the challenges posed by burstiness (chlebus2009maximum, ; chlebus2007stability, ). These algorithms are compared, both analytically and via simulations, against the randomized algorithms, BEB and Quadratic Backoff (QB), where the latter has each successive window size increase as the square of the previous window. The simulation results illustrate the impact of different injection rates, burstiness values, and the number of participants. These findings generally align with the theoretical predictions and, interestingly, the deterministic algorithms compare favorably to BEB and QB, suggesting that randomized approaches may not necessarily be superior. More-recent work by Hradovich et al. (HradovichKK21, ) provides a new average-case analysis of the maximum queue size and latency for several of the same algorithms.
While many deterministic approaches require nodes to have unique IDs, Anantharamu and Chlebus (anantharamu2015broadcasting, ) consider deterministic algorithms where this is not true; furthermore, $n$ is unknown to the algorithm and may not even be finite. In trade, they focus on positive results against leaky bucket adversaries that have the additional constraint that they inject at most one packet per round. An algorithm is activation-based if stations that do not have any packets to send ignore any channel feedback. In terms of positive results, the authors provide a non-adaptive, activation-based algorithm that has bounded latency and is stable for an injection rate $\rho<1/3$ with CD, a non-adaptive full-sensing algorithm that has bounded latency and is stable for $\rho\leq 3/8$ with CD, and an adaptive activation-based algorithm that has bounded latency and is stable for injection rates less than $\rho\leq 1/2$ with no-CD.
In terms of negative results, Anantharamu and Chlebus show that no deterministic algorithm can be fair against an adversary that is able to activate at most $2$ stations in any round—called a $2$-activating adversary—with $\lfloor\rho+b\rfloor=1+b\geq 2$. Additionally, the authors show that no deterministic acknowledgment-based algorithm is fair against a $1$-activating adversary for $\rho<1$ and $b\geq 3-2\rho$. This stands in contrast to the positive results in Chlebus et al. (chlebus2006adversarial, ; chlebus:adversarial, ) for $\rho<1$. Finally, Anantharamu and Chlebus prove that no deterministic algorithm with CD can provide bounded packet latency against a $1$-activating adversary with injection rate at least $3/4$ and $b\geq 1$. It is interesting to compare this result against the algorithm by Chlebus et al. (chlebus2009maximum, ; chlebus2007stability, ) discussed above, which achieves stability, having $O(n^{2}+b)$ total packets queued in the system at any given time (but where there are unique IDs and the number of stations is known).
Anantharamu et al. (anantharamu2009adversarial, ; anantharamu2017adversarial, ) consider deterministic full-sensing and acknowledgment-based algorithms for the case where each participant is subject to an individual injection rate under a window adversary, where the window $w$ is unknown. There are $n$ participants, each with a unique ID in ${1,...,n}$, and $n$ is known. The authors prove a lower bound on latency of $\Omega(w\max\{1,\log_{w}n\})$ for $w\leq n$, and $\Omega(w+n)$ for $w>n$. Upper bounds are established by a full-sensing algorithm with CD, and by an adaptive algorithm without CD, each of which achieve a packet latency of $O(\min(n+w,w\log n))$; Anantharamu et al. also give a full-sensing algorithm without CD that has $O(n+w)$ queue size and $O(nw)$ latency.
In contrast to the above, the remaining results address a model where participants have unique IDs in $\{1,...,n\}$, where $n$ is known. Additionally, participants have access to a global clock. Hradovich et al. (hradovich2020contention, ) consider a problem variant where, in any fixed round, at most $k$ stations may access the multiple access channel; this is called a $k$-restrained channel. The authors design and analyze deterministic algorithms for leaky bucket adversaries with $\rho\leq 1$. The first is a stable adaptive algorithm whose maximum number of queued packets is $O(n^{2}+b)$, which is reminiscent of the bound in (chlebus2009maximum, ). The second is a full-sensing algorithm that is stable for $\rho\leq 1-1/n$. Finally, the authors give an acknowledgment-based algorithm with throughput $\Theta(k/n(\log^{2}n))$, which is complemented by a lower bound that no acknowledgement-based algorithm can achieve a throughput better than $\min\{k/n,1/(3\log n)\}$. Simulations indicate that the proposed algorithms perform favorably, with respect to queue size and number of channel accesses, relative to well-known backoff algorithms.
Bienkowski et al. (bienkowski:distributed, ; BienkowskiJKK12, ) consider a model where each participant has a queue, and messages may include control bits. The adversary considered here differs from the leaky-bucket and window adversaries discussed earlier: at any time, an arbitrary number of packets can be injected into the queue of any participant. The metrics of interest are the total packets in all queues, called total load, and the maximum queue size. The authors prove that their deterministic algorithm achieves a total load that is an additive $O(n^{2})$ amount larger than the best offline algorithm, and the maximum queue size is at most an $n$-factor larger.
In work by Garncarek et al. (garncarek2018local, ), participants receive no channel feedback, and they may inform their decisions using knowledge of their own local queue, or information gained about other queues. In particular, an adaptive algorithm may store bits in memory that reflect queue histories, while non-adaptive (memoryless) algorithms may not. Both types of algorithms can check whether or not their queue is empty. The authors design a deterministic, adaptive algorithm that guarantees that all stations have bounded queue size $L_{\mbox{\tiny max}}$ for any $\rho$ and $b$; that is, the algorithm is universally stable. The amount of state that each participant maintains is at most $n\log(L_{\mbox{\tiny max}})$ bits. The authors also prove that there is no non-adaptive algorithm that is stable for an injection rate of $O(1/\log n)$, and they give a non-adaptive algorithm that is stable for an injection rate of $d/\log^{2}n$, for some constant $d>0$.
More recently, Garncarek et al. (garncarek2019stable, ) operate in the same model, with a focus on algorithms that keep very little or no information about the previous state of the system. Here, the authors show a deterministic algorithm that is stable for $\rho=\Omega(1/\log n)$ given that only $O(1)$ bits are used as memory to store some history of events, and each participant can check its queue size (not just whether it is empty). Stability can also be achieved when the rate is increased to a constant, and with zero bits of memory; however, $b$ must be upper bounded by some value known a priori.
Summary. There are several significant messages conveyed by the literature surveyed above. Randomized backoff algorithms continue to disappoint in that they are stable only for sub-constant injection rates. Interestingly, in terms of latency, deterministic algorithms can achieve performance that rivals and even surpasses traditional randomized backoff algorithms. However, most of these results make assumptions about unique IDs and/or a priori knowledge of $n$. The exception is the algorithm by Anantharamu and Clebus (anantharamu2015broadcasting, ), which assumes a $1$-activating adversary. This leads to a more general message: while stability and fairness seem like natural qualities that we desire in algorithm design, it is often not possible to achieve both simultaneously when there is even a little (i.e., constant) burstiness. Therefore, all of this paints a fairly dismal picture of what is possible under window and leaky bucket adversaries, especially for deterministic algorithms, given the many impossibility results that apply.
What avenues are there for positive results? As we will discuss later in Section 3.3, new randomized approaches can offer better performance. However, fundamentally, perhaps the AQT model is overly pessimistic; this has been investigated in more general settings (e.g., (BKS14, ).) Here, focusing on whether queue sizes can become unbounded under a very specific injection pattern, and under an infinite stream of packets, may lead to results that are artificially bleak. In the next section, we turn our attention to work that is concerned with a finite stream of packets that arrive under an arbitrary schedule.
3.2. Deterministic Approaches
  • “Anyone who attempts to generate random numbers by deterministic means is, of course, living in a state of sin.”   – John von Neumann
In this section, we depart from parameterized adversaries and notions of stability, while still focusing on deterministic contention resolution algorithms. Here, our performance metric is latency (recall Section 1.1.3). Much of the contention-resolution work in this setting is referred to as $k$-selection, which is described as follows. There are $n$ participants of which $k$, for $1\leq k\leq n$, activate over time, each with a single packet to transmit. The activation times are selected by an adversary. As before, most of the literature operates in a model where each participant has a unique ID. Furthermore, $n$ is typically assumed to be known; throughout this section, we assume this to be true unless explicitly stated otherwise. The goal is to minimize the latency (recall from Section 1.1 ) required for all $k$ participants to successfully send their respective packets.
Figure 5. Depiction of a transmission matrix presented in the work by De Marco and Kowalski (doi:10.1137/140982763, ). For two participants, $u$ and $w$, we illustrate the intervals over which they respectively may transmit on the shared channel.
Developing Intuition.
Many results put a spin on John von Neumann’s quote by generating deterministic results from randomized approaches. Specifically, a common technique in this area is to design a randomized algorithm, and then use the probabilistic method (alon2016probabilistic, ) to argue that a deterministic analog exists. Here, we focus on the result by De Marco and Kowalski (doi:10.1137/140982763, ); however, related approaches are used in several other papers to give existence proofs for efficient transmission schedules (DBLP:conf/icdcs/MarcoKS19, ; de2013contention, ; DEMARCO20171, ; Chlebus:2016:SWM:2882263.2882514, ; ChlebusK04, ; ChlebusGKR05, ).
De Marco and Kowalski (doi:10.1137/140982763, ) consider the dynamic $k$-selection problem where there is a global clock, $k$ is unknown, and no-CD. Their main result is a non-adaptive, deterministic algorithm for contention resolution with latency $O(k\log(n)\log\log(n))$. Notably, this result almost matches a previous lower bound by Greenberg and Winograd (greenberg1985lower, ) in $\Omega(k\log(n)/\log k)$ for adaptive algorithms under the static case with collision detection.
A key component of the authors’ (doi:10.1137/140982763, ) approach is a transmission matrix, $\mathcal{M}$, consisting of $\lg n$ rows and $\ell=\Theta(n\log n\log\log n)$ columns. The way in which participants use $\mathcal{M}$ to transmit on the channel is as follows. Let $\mathcal{M}_{i,j}$ denote the entry at row $i$ and column $j$; each such entry is a list of IDs. Upon activating, in each slot participant $u$ scans left-to-right in row $1$ of $\mathcal{M}$. In this row, there is a designated interval of slots $\mathcal{I}_{1,u}$ for $u$; specifically, $u$ rounds up the current (global) slot index to the nearest multiple of $2\lg\lg n$ and this is the start of $\mathcal{I}_{1,u}$, which has length $\Theta(\log n\log\log n)$. If $u$ belongs to the list at $\mathcal{M}_{1,j}$ for $j\in\mathcal{I}_{1,u}$, then $u$ sends its packet in slot $j$. If at any point $u$ succeeds in $\mathcal{I}_{1,u}$, then $u$ terminates; otherwise, $u$ immediately proceeds to row $2$ and starts executing as prescribed by $\mathcal{I}_{2,u}$. This process through $\mathcal{M}$ continues until all rows are exhausted. The length of interval, $\mathcal{I}_{i,u}$, in each row $i$ is $\Theta(2^{i}\log n\log\log n)$. Figure 5 illustrates a transmission matrix.
How is membership in $\mathcal{M}_{i,j}$ determined? This is decided randomly: participant $u$ belongs to $\mathcal{M}_{i,j}$ with probability approximately $\frac{1}{2^{i}2^{j\mod\lg\lg n}}$, for $1\leq i\leq\lg n$ and $j\geq 1$. Note that the first term of the sending probability (the $1/2^{i}$ term) and also the interval length (specifically, the $2^{i}$ term) are reminiscent of the $i^{\mbox{\tiny th}}$ window for the Bernoulli version of BEB (recall Sections 2.1 and 3.1). Consequently, we should not be surprised at a $\Omega(k\log n)$ latency.
How does contention behave under $\mathcal{M}$? For illustrative purposes, consider $\Theta(k)$ participants executing the schedule in rows $1$ and $\lg k$. In row $1$, an active participant $u$ may contribute significantly to contention, but only for its small interval $\mathcal{I}_{1,u}$. For example, if $\Theta(k)$ participants all activate simultaneously (i.e., a batch), then their contribution to contention is $\Theta(k)$ and no participant succeeds; however, this occurs only over $\Theta(\log n\log\log n)$ slots in row $1$. Conversely, the $\Theta(k)$ participants may be spread out over row $1$, each contributing only $O(1)$ contention in their own respective, disjoint interval of $\Theta(\log n\log\log n)$ slots; we will return to these two extremes momentarily.
In row $\lg k$, the sending probability ranges between $\Theta(1/k)$ and $\Theta(1/(k\log n))$. Thus, a batch of $\Theta(k)$ participants make a more muted contribution to contention, ranging from $\Theta(1)$ to $\Theta(1/\log n)$ over a much larger interval $\mathcal{I}_{\lg(k)}$ consisting of $\Theta(k\log n\log\log n)$ slots. This range of contention is “good” in the sense that, recalling Equation 1, in the absence of other major contributions to contention, we expect participants to succeed. Is this the case?
Recall the contribution to contention from participants in row $1$ using the two qualitatively different cases discussed above: their contribution may be high for a “small” number—specifically, $\Theta(\log n\log\log n)$—slots of $\mathcal{I}_{\lg(k)}$, or only $O(1)$ over perhaps all of $\mathcal{I}_{\lg(k)}$. Therefore, accounting for these two cases, in row 1 there should be many disjoint periods of $\Theta(\log\log n)$ contiguous slots where contention is good. More generally, given the contribution to contention from other rows, the authors show that, w.h.p. in $n$, an important property holds: in every interval of length $\Theta(k\log(n)\log\log(n))$ slots, $O(k/\log n)$ participants proceed from row $\lg(k)$ to row $\lg(k)+1$ of the transmission matrix. De Marco and Kowalski (doi:10.1137/140982763, ) show that a random transmission matrix yields the above property with strictly positive probability. Therefore, by the probabilistic method, the authors argue that there must exist a matrix with this property.
But what about these $O(k/\log n)$ participants that keep executing in row $\lg(k)+1$? Interleaved with the execution is the deterministic result by Komlós and Greenberg (komlos:asymptotically, ) (recall Section 2.2), which addresses the static case. Specifically, given a batch of $m$ out of $n$ participants with a packet to send, where both $m$ and $n$ are known, all $m$ participants succeed in time $O(m+m\log(n/m))$. Here, in the dynamic $k$-selection problem, $m$ is not known, and so the result by Komlós and Greenberg is executed for values of $m=1,2,...,\lg n$, which contributes an extra logarithmic-in-$n$ factor to the run time. Thus, the $m=O(k/\log n)$ participants that remain after row $k$ succeed within an additional $O(m+m\log(n/m))\log n$, which is $O(k\log(n\log n))$ by substituting in for $m$.
Finally, note that by the sum of a geometric series, the number of slots in $\mathcal{M}$ over which any participant executes prior to reaching row $\lg(k)+1$ is $O(k\log n\log\log n)$. Adding to this the (asymptotically smaller) $O(k\log(n\log n))$ slots for the Komlós and Greenberg result yields a total latency of $O(k\log n\log\log n)$.
3.2.1. Deterministic $k$-Selection
Greenberg and Winograd (greenberg1985lower, ) consider a multiple access channel with CD, for the two cases where $k$ is known and unknown. The authors prove a lower bound for deterministic algorithms of $\Omega(k\log n/\log k)$ in the static case, which of course applies to the dynamic case. This is contrasted against the upper bounds, claiming that the deterministic algorithms in Capetanakis (capetanakis:generalized, ), Hayes (hayes:adaptive, ), and Tsybakov and Mikhailov (tsybakov1978free-russian, ; tsybakov1978free, ) provide a nearly-tight latency of $\Theta(k+k\log(n/k))$.
With regard to upper bounds, De Marco and Kowalski (doi:10.1137/140982763, ) consider the case with a global clock and no-CD, as discussed above. Subsequently, De Marco et al. (DBLP:conf/icdcs/MarcoKS19, ) examine the power of acknowledgement-based versus non-acknowledgement-based deterministic non-adaptive approaches when there is no global clock and under the no-CD setting. For unknown $k$, they give a non-acknowledgement-based algorithm with latency $O(k^{2}\log n)$, and show that this can be improved to $O(k^{2}\log n/\log k)$ with an acknowledgment-based algorithm. These results are (nearly) tight; the authors also provide lower bounds on latency of $\Omega(k^{2}/\log k)$ for both cases. This illustrates a startling contrast with the case for known $k$, where the authors give an acknowledgement-based algorithm with $O(k\log k\log n)$ latency, which is known to be nearly tight (ClementiMS03, ).
3.2.2. Deterministic Wakeup
[…]
Continuing in a model with no-CD, but with a global clock, De Marco and Kowalski (DEMARCO20171, ; de2013contention, ) consider non-adaptive solutions. Here, there is a parameter $k$, for $1\leq k\leq n$, which is an upper bound on the number of stations that may become active over time. Of these stations, $s$ is the time (measured from the first slot) at which the first station becomes active. The authors look at three settings: where $s$ is known and $k$ is unknown, when $k$ is known and $s$ is unknown, and when both $k$ and $s$ are unknown. In the first two settings, the authors give algorithms that achieve a latency of $O(k\log(n/k)+1)$, while their solution for the third scenario has a latency of $O(k\log(n)\log\log(n))$. A lower bound of $\Omega(k\log(n/k)+1)$ is given, which shows that their first two upper bounds are asymptotically optimal, and the third is within a $O(\log\log n)$-factor of optimal.
Chlebus et al. (Chlebus:2016:SWM:2882263.2882514, ) consider a multi-channel scenario with no-CD and no global clock. Here, all stations may access all $b\geq 1$ channels at once, and with the same parameter $k\geq 1$ being an upper bound on the number of stations that become active over time. The authors give a non-adaptive, deterministic algorithm for the cases where $k$ is unknown that has latency $O(k\log^{1/b}k\log n)$. When $b=\Omega(\log\log n)$, a faster deterministic algorithm is possible, with delay $O((k/b)\log n\log(b\log n))$, and a lower bound of $\Omega((k/b)\log(n/k))$ is shown for deterministic algorithms. For ease of presentation, we note that the authors also consider jamming (a topic discussed further in Section 4), where each channel is jammed independently and uniformly with known probability $p$ per round, where $10<p<$ 1. Here, Chlebus et al. (Chlebus:2016:SWM:2882263.2882514, ) give two deterministic algorithms that succeed with probability $1-O(1/\texttt{poly}(n))$, the first with latency $O(k\log(n)\log^{1/b}k/\log(1/p))$, and the second with latency $O((k/b)\log(n)\log(b\log(n))/\log(1/p))$ when $b=\Omega(\log(b\log n))$.
[…]
Finally, under the SINR channel model, Jurdziński et al. (jurdzinski2013distributed, ) give a deterministic algorithm with $O(\log^{2}n)$ makespan when there is a global clock; however, without a global clock, the authors show a $\Omega(n)$ makespan. Complementing this result, Jurdziński and Stachowiak (jurdzinski2015cost, ) show that any algorithm with access to a global clock must incur $\Omega(\log^{2}n)$ makespan in the classical (non-SINR) channel model.
Summary. Much of the theoretical territory has been mapped out with some nearly-matching upper and lower asymptotic bounds on latency, and the results are less bleak than those surveyed under the AQT model. Notably, with no-CD and assuming a global clock, we see this for both $k$-selection (De Marco and Kowalski (doi:10.1137/140982763, )) and for wakeup (De Marco and Kowalski (DEMARCO20171, ; de2013contention, )), where the latter result depends on some limited system information.
Despite providing a rosier picture, deterministic approaches for contention resolution fall short of being able to offer asymptotically-optimal latency in the dynamic case; this is unsurprising given the analogous situation in the static case (Section 2.2). Similarly, deterministic solutions to the wakeup problem yield high latency in cases without a global clock, or where key parameters are unknown. A natural place to look for improvement—perhaps even asymptotic optimality—is the domain of randomized algorithms, which we survey next.
3.3. Randomized Approaches for Contention Resolution
As we saw in Section 2, randomized backoff algorithms provide asymptotically-optimal makespan for the static case without requiring unique IDs, CD, or knowledge of $n$. Thus, there is reason to hope that randomized approaches can also provide strong results in the dynamic case.
Developing Intuition.
A common theme in the design of randomized algorithms for contention resolution is to decompose the dynamic case into a sequence of static cases. Here, we describe one method by which this decomposition is accomplished (for examples, see (BenderKPY18, ; bender:contention, ; bender2020contention, )).
When a packet $u$ arrives, it listens to the channel to determine whether it is already in use by other packets. Any packets that have arrived ahead of $u$ may still be attempting to succeed and, if so, they should also be emitting a “busy signal”, which we describe momentarily. If $u$ detects this busy signal, then it listens until this signal ceases, at which point it can start attempting to transmit. In this way, there are always two groups of packets: (1) active packets that are trying to transmit, while maintaining a busy signal, and (2) idle packets that are waiting for their turn on the channel. Given this isolation, active packets can execute a contention-resolution algorithm for the static case in order to succeed; afterwards, any idle packets can become active and repeat this process.
The busy signal can take different forms. For example, given CD and a global clock, active packets can emit a ”busy” signal in each even-numbered slot; see Figure 6. Any idle packet that hears this signal or a collision infers that active packets are present and keeps waiting. However, having all packets send in every second slot may not be desirable, since sending consumes energy (SINHA201714, ; feeney2001investigating, ; polastre:telos, ; wang2006realistic, ). To reduce this cost, packets may emit the busy signal probabilistically, which can greatly reduce the sending cost; however, there is now a probability of error in the situation where some packets may still be active, but fail to emit a busy signal. When this happens, any idle packets will become active prematurely and interfere with the already-active packets.
This interference is mitigated differently, depending on the setting. Bender et al. (BenderKPY18, ; bender:contention, ) assume CD and $O(\log n)$ control bits (recall Section 1.1.4) in each transmission on the shared channel. The busy signal occurs probabilistically every $O(1)$ slots, with each active packet sending an expected $O(1)$ times to maintain the signal. Interleaved with the busy signal are slots used to estimate the number of active packets, $m\leq n$, which allows them to execute TSTB (recall Section 2.1) starting with a window of size $\Theta(m)$. Additionally, a leader is elected from the set of active packets to broadcast how many slots remain in the execution of TSTB; this serves as a busy signal during the execution of TSTB. The leader is frequently changed so that each packet is responsible for $O(1)$ transmissions in expectation. Newly-arrived idle packets need only listen for a handful of slots to receive this information, which allows them to (i) avoid becoming active prematurely and (ii) avoid any further listening until the current execution of TSTB is finished. While the above process can fail—for example, the estimate is poor, the busy signal fails, or active packets do not succeed in the execution of TSTB—this occurs with probability at most polynomially small in the number of active packets, which allows for $O(n)$ expected latency. Additionally, the expected number of channel accesses performed per packet is $O(\log\log^{*}n)$. This low number of accesses translates is important to modern networks where devices are battery-powered, given that listening and sending on the channel consumes limited energy.
We note that while a busy signal has been used to achieve strong algorithmic results, it may be problematic in practice, since a constant fraction of the active slots are wasted on this synchronization.
Figure 6. An example of how a busy signal may be used to decompose the dynamic case into a series of static cases. Here, three packets (colored red) arrive in slot 1 and issue the busy signal in slots 2, 4, and 6. During this time, other packets (colored blue) arrive, but remain idle, since they detect the busy signal. Once all red packets succeed, slot 8 is empty, and the blue packets become active.
3.3.1. Randomized $k$-Selection
We extend our discussion above by continuing to survey the results that employ randomized algorithms for the dynamic $k$-selection problem. In the context of randomized algorithms, results focus the case where $k=n$. We note that the jamming-resistant contention-resolution results (presented next in Section 4) clearly apply when there is no jamming, although we defer their presentation until later.
De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ) explore randomized solutions with no-CD. When $n$ is known (or a known to within a constant multiplicative factor), they provide a non-adaptive algorithm with $O(n)$ latency, which is asymptotically optimal. For the case where $n$ is unknown, a lower bound of $\Omega(n\log n/(\log\log n)^{2})$ latency is shown to hold for any non-adaptive, randomized algorithm; thus, there is an asymptotic separation in latency between adaptive and non-adaptive randomized algorithms. Additionally, for $n$ unknown, the authors provide a strongly non-adaptive, non-acknowledgment-based algorithm with $O(n\log^{2}n)$ latency; this latency can be improved to $O(n\log n/(\log\log n)^{2})$ with an acknowledgment-based algorithm, which comes close to matching their lower bound. Lastly, when $n$ is unknown, the authors provide an adaptive algorithm that achieves $O(n)$ latency. Informally, this algorithm employs LE to break the dynamic problem into a sequence of static contention-resolution problems, as discussed earlier in this section.
A more recent result by Bender et al. (bender2020contention, ) shows that w.h.p. constant throughput can be achieved by an adaptive algorithm with no-CD. In contrast, to the adaptive algorithm (for unknown $n$) of De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ), the result by Bender et al. (bender2020contention, ) allows packets to terminate immediately upon success.
When CD is available, Bender et al. (bender:contention, ) achieve $O(n)$ latency using expected $O(\log\log^{*}n)$ channel accesses per packet; this is shown to be asymptotically optimal by Chang et al. (ChangJP19, ). As discussed in Section 1.1.3, this low number of channel accesses translates is important to modern networks where devices are battery-powered, given that listening and sending on the channel consumes energy. We highlight that LE plays a key role in keeping energy costs low.
In a multi-channel setting, with and without CD, Christodoulou et al. (christodoulou2018strategic, ) consider a game theoretic setting where participants may deviate from any prescribed protocol. The authors give protocols from which no participant will unilaterally deviate and which provide probabilistic latency guarantees.
Finally, recent work by Bender et al. (bender:coded, ) investigates a problem variant where at most an integer $\kappa\geq 1$ simultaneous successful transmissions are possible. This model is reminiscent of (hradovich2020contention, ) (in Section 3.1), and it captures the newer radio technology that can decode a limited number of concurrent transmissions. Here, the notion of CD alters slightly, where participants can distinguish between an empty slot and a slot with multiple transmissions; however, participants do not learn whether the number of transmissions in any slot exceeds or is at most $\kappa$ (unless they are one of the senders). Here, the authors show that a throughput of $1-o(1)$ is possible.
3.3.2. The Wakeup Problem
There is a known separation between static and dynamic wakeup. Specifically, recall that for static wakeup with no-CD, Kushilevitz and Mansour (kushilevitz1998omega, ) show a lower bound of $\Omega(\log n)$, and an asymptotically matching upper bound is achieved via a randomized algorithm by Bar-Yehuda et al. (bar1992time, ). However, for dynamic wakeup, Jurdzinski and Stachowiak (jurdzinski2015cost, ) prove a lower bound of $\Omega(\log^{2}n)$ in the case of no-CD.
Under no-CD, Jurdziński and Stachowiak (jurdzinski2005probabilistic, ; jurdzinski2002probabilistic, ) present randomized algorithms with probability of error $\epsilon$ for solving the wakeup problem under different settings: where participants are labeled or unlabeled, where there is a global clock or not, and where $n$ is known or unknown. We note that, while deterministic approaches rely on the use of labels, this is an example of labels being used with a randomization solution; in this case, the labels are used to approximate system size. The authors are concerned with uniform algorithms (recall Section 2.3). Perhaps unsurprisingly, aspects of these wakeup algorithms are reminiscent of exponential backoff. One of the main findings is a near-exponential separation between the global and local clock settings when $n$ is unknown and there no labeling; the former can be solved with latency $O(\log(n)\log(1/\epsilon))$, while the latter requires $\Omega(n/\log n)$ latency. In the cases where $n$ is known or the participants are labeled (or both), the run time is $O(\log n\log(1/\epsilon))$. For uniform algorithms, the authors also show a nearly-tight lower bound of $\Omega\left(\frac{\log(n)\log(1/\epsilon)}{\log\log(n)+\log\log(1/\epsilon)}\right)$ in the case where $n$ is known and the participants are labeled. As stated earlier in Section 2.3, this bound improved by Farach-Colton et al. (farach-colton:lower-bounds, ) to $\Omega(\log(n)\log(1/\epsilon))$, and then later shown to hold for all randomized algorithms by Newport (newport2014radio, ; newport:radio-journal, ).
Under the SINR model (recall Section 2.3), and in the absence of a global clock and assuming participants know a polynomial upper bound on the system size, Jurdziński and Stachowiak (jurdzinski2015cost, ) give a Monte Carlo algorithm that has w.h.p. $O(\log^{2}n/\log\log n)$ latency, and a Las Vegas algorithm with expected $O(\log^{2}n/\log\log n)$ latency. We note that this is better than the deterministic algorithm in (jurdzinski2013distributed, ) of $O(\log^{2}n)$ latency (recall Section 3.2.2).
For a single-hop radio network, Chlebus et al. (Chlebus:2016:SWM:2882263.2882514, ) consider $k$ devices becoming active over time, with access to a known number of $b$ channels. The model used is no-CD and no global clock. Here, all stations may access all channels at once. When $k$ is known, Chlebus et al. present a randomized algorithm with error $\epsilon$ that has latency $O(k^{1/b}\ln(1/\epsilon))$.
[…]
Summary. For the $k$-selection problem, a key message is that we can get asymptotically optimal $O(k)$ latency. For the wakeup problem, the situation is also significantly improved by randomization. Specifically, compared to the results surveyed in Section 3.2.2, we see that randomization often allows for $O(\texttt{polylog}(n))$ latency in contrast to $\omega(n)$ latency in the single-hop scenario; however, in the local clock setting without knowledge of key parameters, the latency remains high. That said, in comparison to the results from the AQT and deterministic approaches, it is clear that randomization can offer significant improvements.
Another aspect of randomized approaches to the selection problem is the focus on the number of access attempts, which translates to energy consumption. The $O(\log\log^{*}n)$ complexity achieved by Bender et al. (bender:contention, ) exploits the ability to send and receive messages of $\Theta(\log n)$ bits. Under the more standard ternary feedback model, the best results achieve a polylogarithmic in $n$ number of access attempts.
Finally, we note that while many of the randomized results for contention resolution achieve asymptotically-optimal latency, they are fragile in the sense that even a small amount of channel interference can cause the corresponding algorithm to fail. For example, the result of Bender et al. (bender:contention, ) estimates the number of packets contending for channel access; even a single slot with interference can cause a gross overestimate. Another instance is a result by De Marco and Stachowiak (DeMarco:2017:ASC:3087801.3087831, ), which depends on coordination messages issued by a leader; such messages can be disrupted by interference. The algorithm by Bender et al. (bender2020contention, ) is also unable to maintain asymptotically-optimal latency in the face of jamming. Thus, robustness to malicious interference presents a challenge, and it is the subject we turn to next.
4. Jamming-Resistant Contention Resolution
  • “I bought the jammer from a local electrical store and asked the police if it was OK and they said it was - it’s great as it has stopped the problem, but some of the local shops aren’t happy.”
                          – Father Michele Madonna of Santa Maria di Montesanto Church (priest, )
A jamming attack is characterized by deliberate collisions on the multiple access channel. By deviating from a prescribed protocol, a jamming device (jammer) can launch a denial-of-service attack, preventing other devices from being able to communicate (DBLP:reference/algo/RichaS16, ). There have been several well-publicized instances of jamming (FCC:humphreys, ; marriott, ; priest, ; nicholl, ), behavior that may become more common given the trend towards wireless personal devices. For example, alluding to the quote above, Father Madonna installed a store-bought jamming device in his church to prevent attendees from disrupting his services with cellphone usage. However, this device also interfered with nearby wireless devices, with shop owners complaining that credit card payment systems were affected.
There are a number of papers addressing applied security considerations with respect to jamming (for examples, see (bayraktaroglu:performance, ; xu:feasibility, ; deejam:wood, ; aschenbruck:simulative, ; wood:denial, ; xu:jamming, ) and surveys (mpitziopoulos:survey, ; VADLAMANI201676, )). For example, spread spectrum techniques (spread:foiling, ; navda:using, ), rerouting defenses (wood2003jam, ; han2018hierarchical, ), and specialized antenna technology (zheng2018design, ; mcmilin2015gps, ). From a theory perspective, several fundamental distributed computing challenges—such as broadcast (koo2, ; gilbert:near, ; gilbert:making, ; chen:broadcasting, ), leader election (gilbert:malicious, ; gilbert2009malicious, ), gossip (gilbert:interference, ; dolev:gossiping, ), node discovery (meier:speed, ), and simple point-to-point communication (king:conflict, ; DBLP:journals/dc/KingPSY18, )—can be solved on a shared channel despite jamming (see (young:overcoming, ; Bender:2015:RA:2818936.2818949, )). Here, we focus on prior contention-resolution results that offer provable guarantees on throughput or latency, despite jamming.
Developing Intuition.
A jamming adversary poses at least two significant new challenges over the dynamic-arrivals model considered in Section 3. First, it may not be obvious that non-trivial makespan guarantees are possible; for example, if jamming occurs in every slot in perpetuity, then no participants can succeed. Therefore, we must either constrain the jamming power of the adversary (e.g., by assuming a bounded adversary), or alter our measure of throughput to address this case (e.g., such as notions of competitiveness and implicit throughput). Second, the algorithms surveyed so far that rely on channel feedback to adjust contention—for example, by estimating $n$ (such as in (BenderKPY18, ; DBLP:journals/jacm/GreenbergFL87, )) or by exchanging messages between participants—are vulnerable, since jamming allows the adversary to create collisions and prevent communication.
With regard to the first challenge, a common approach is to measure performance over non-jammed slots; for example, this is done by Chang et al. (ChangJP19, ) and Bender et al. (bender:how, ). Specifically, assuming that at least one participant is present in the system, a slot that is not jammed, but where no participant succeeds, is said to be wasted. Guaranteeing that only a small constant fraction of slots are wasted allows us to derive meaningful results despite an adversary that jams aggressively, perhaps even in every slot.
Addressing the second challenge of adjusting the contention requires care, and this is demonstrated by the elegant algorithm by Chang et al. (ChangJP19, ). Participants follow some simple rules for updating the sending probability in each slot: increasing if there is silence, decreasing if there is noise, and leaving it unchanged in the case of a success. The adversary may lower the sending probabilities of participants by jamming (i.e., adding noise), thus depressing contention. However, the update rule allows for contention to quickly recover once the jamming ceases (although, it may resume again later), and few slots are wasted as a result. Notably, this design allows the algorithm to achieve roughly $1/e$ throughput in the non-jammed slots, which rivals that of slotted ALOHA, despite the more-challenging adversarial setting.
Survey of the Literature.
Broadly speaking, the jamming adversaries treated in the literature on contention resolution can be categorized as bounded or unbounded. The former has constraints that limit the amount of jamming; typically, this is specified by a jamming rate or by a fraction of the slots that can be jammed within any contiguous set of slots. In contrast, as the name suggests, an unbounded adversary may jam arbitrarily.
An early series of results against a bounded adversary was initiated by Awerbuch et al. (awerbuch:jamming, ; DBLP:journals/talg/AwerbuchRSSZ14, ), who consider the following model. There are $n$ honest participants equipped with CD, and the network is saturated (recall Section 2.1). For some $T\in\mathbb{N}$ and constant $10<\epsilon<$ 1, there is an adversary that can jam at most $(1-\epsilon)w$ of the slots in any consecutive $w\geq T$ slots. This adversary is referred to as $(T,1-\epsilon)$-bounded, and the value of $\epsilon$ is unknown to participants.
Such bounded jamming may be well-motivated in practice, where aggressive jamming can reveal that an attack is occurring, as well as the location of the malicious device(s) (for examples, see (liu2009localizing, ; liu2013error, ; xu:feasibility, )), or correspond to a limited power supply. The adversary is also adaptive, meaning that the adversary knows the full history of the actions performed by honest nodes for all the previous slots, and may use this information to inform its jamming strategy in the current slot; however, the adversary does not know what actions participants are taking in the current slot. When the adversary jams a slot, all listening participants hear jamming; this is referred to as a 1-uniform adversary.
Awerbuch et al. (awerbuch:jamming, ; DBLP:journals/talg/AwerbuchRSSZ14, ) define an algorithm to be constant competitive if, for any sufficiently large number of slots, the honest nodes perform successful message transmissions in a constant fraction of the non-jammed steps. The authors provide a randomized algorithm that is constant competitive with high probability (w.h.p.) in $n$, when all participants know a common parameter, $\gamma$$=O(1/(\log T+\log\log n))$, and the algorithm is run for a sufficiently large number of slots. Roughly speaking, in a slot, the algorithm dictates that each participant $u$ sends a message with probability $p_{u}$. If $u$ does not send, it listens and updates $p_{u}$ via a multiplicative factor which can increase or decrease, depending on whether or not the slot is idle. Notably, this multiplicative weights update method shares common ground with the later algorithm by Chang et al. (ChangJP19, ) discussed above.
Subsequently, this line of work has been extended to several challenging settings, showing constant competitiveness so long as $\gamma$ is known to all the participants, and the algorithm is executed for a sufficiently large number of slots. Richa et al. (richa:jamming2, ) consider a multi-hop network modeled as a unit disk graph (UDG). Informally, under the UDG model, each participant can broadcast within a disk centered about its location, and two participants may communicate if their respective disks intersect (see Schmid and Wattenhofer (schmid2006algorithmic, ) for discussion of the UDG model). First, Richa et al. (richa:jamming2, ; DBLP:journals/dc/RichaSSZ13, ) consider the setting where the adversary is 1-uniform and the UDG is connected, or each participant has at least $2/\epsilon$ neighbors. Then, against a $1$-uniform $(T,1-\epsilon)$-bounded adversary, the authors provide an algorithm, named Jade, that is w.h.p. constant competitive. Second, Richa et al. (richa:jamming2, ) prove that, if the graph is not connected, or the adversary is $2$-uniform—for any slot, the adversary may choose any partition of the $n$ participants and have the members of one partition hear jamming, while the others do not—and participants each have $o(1/\epsilon)$ neighbors, then Jade can fail to be constant competitive. Against a reactive jamming adversary—who can learn whether or not the channel is idle in the current slot and can choose to jam based on this information—Richa et al. (richa:jamming3, ; DBLP:journals/ton/RichaSSZ13, ) show that constant competitiveness is still possible.
Under the SINR channel model (recall Section 2.3), Ogierman et al. (ogierman:competitive, ; ogierman2018sade, ) address jamming in a more-general network topology than prior results on competitiveness. The authors provide a multiplicative weights update algorithm that is $2^{-O(1/\epsilon)^{2/(\alpha-2)}}$-competitive, where $\alpha>2$ is a parameter of the SINR model known as the path-loss exponent, which corresponds to the degree of signal attenuation. This competitiveness holds if the algorithm is executed for a sufficiently large number of slots against a $1$-uniform adversary and the transmission range contains at least one node, or at least $2/\epsilon$ nodes within the transmission range of every node.
In the case where multiple networks inhabit the same terrain, bandwidth may be scarce and inter-network communication may not be possible. In this setting, Richa et al. (richa:jamming4, ) present a multiplicative weights update algorithm that allows each of the $K$ networks to obtain $\Omega(\epsilon^{2}\min\{\epsilon,1/\mbox{poly}(K)\})$-fraction of the non-jammed slots.
A different approach is taken by Bender et al. (bender:how, ; BenderFGY19, ), who address an adversary that is not constrained in its jamming. A packet starts in an “inactive” state by listening to the channel, and if a sufficient number of empty slots are heard, the packet becomes “active” in the sense of making transmission attempts. This “throttling” mechanism gives time for contention to decrease, thus allowing active packets to succeed. An active packet $u$ will make transmission attempts, but in contrast to updating based on feedback in each slot, $u$ monotonically reduces its sending probability as roughly $1/s_{u}$, where $s_{u}$ is the number of slots since $u$ became active (i.e., the age of $u$). Packet $u$ also listens (when not transmitting), and if a sufficient number of empty slots are heard, it resets $s_{u}=1$. Here, they show expected constant throughput, with each packet making an expected number of transmission attempts that is polylogarithmic in $n+J$, where $J$ is the total number of jammed slots during which there exists at least one packet.
Recently, Bender et al. (bender:fully, ) showed how to achieve $\Omega(1)$ implicit throughput (recall Section 1.1.3) with $O({\texttt{poly}({\log n})})$ channel accesses (both sending and listening), despite an adaptive jamming adversary. This result applies in the channel setting with CD (i.e., ternary channel feedback), but with no control bits allowed, unlike the energy-efficient result in (BenderKPY18, ). The algorithm, named Low-Sensing Backoff, is based on multiplicative weights updates, and in that regard, it shares common ground with some prior results (such as (ChangJP19, ; richa:jamming2, ; richa:jamming3, ; richa:jamming4, ; awerbuch:jamming, ; ogierman2018sade, )). However, in comparison to these results, Low-Sensing Backoff greatly limits the frequency at which packets can sense the channel in order to update, while the lack of control bits greatly limits the amount of information they receive in each update. Analogous results are given for the cases where the number of packets is infinite, and where the adversary is reactive.
The AQT setting (recall Section 3.1) has been extended to incorporate jamming. Anantharamu et al. (anantharamu2019packet, ; anantharamu2011medium, ) extends a leaky-bucket adversary to ${(\lambda,\rho,T)}$-bounded leaky bucket adversary who can also jam $\lambda t+b$ slots in any window of consecutive $t$ slots; here, $\lambda$ is the jamming rate, while $b$ is the burstiness parameter. Deterministic algorithms are presented, with and without CD, giving a range of bounds on queue size and latency parameterized by the injection rate, burstiness, and the known number of stations $n$, each of which has a unique ID in the range $[0,...,n-1]$.
In a related setting, but with a single sender and receiver, Anta et al. (DBLP:journals/tcs/AntaGKZ17, ) consider an adversary whose jamming capability is parameterized by a rate, $\rho\in[0,1]$, and the length of the largest burst of jammed packets, $\sigma\geq 1$; both parameters are known to the algorithm. Each packet has a fixed-length header, along with a payload whose size may be adjusted by the sender. The authors derive a lower bound on the optimal rate of transmission for payload information (i.e., goodput), and also develop an adaptive deterministic algorithm that is optimal.
Jurdziński et al. (DBLP:conf/waoa/JurdzinskiKL14, ) also consider a single sender and receiver, but in contrast to a parameterized adversary, they consider: an unbounded adversary, and packets of different sizes may arrive over time. If the adversary jams at any time during which a packet is transmitted, the transmission fails and the sender learns of this immediately. The authors derive results on the ratio between the throughput achieved by the algorithm and the throughput achieved by the best scheduling under identical arrival and error patterns; this is referred to as relative throughput. The authors give a deterministic online algorithm that achieves optimal relative throughput.
An impossibility result is given by Bender et al. (bender2020contention, ), who show that with no-CD, no algorithm can guarantee a packet success in time linear in the system size with better than constant probability in the presence of jamming; this result holds even if against a non-adaptive adversary. Subsequently, Chen et al. (DBLP:conf/podc/ChenJZ21, ) give an algorithm in the no-CD case that achieves a trade-off between the throughput and the amount of jamming tolerated. In the challenging case where a constant fraction of slots are jammed, $O(1/\log n)$ throughput is shown, which aligns with the lower bound of Bender et al. (bender2020contention, ).
Summary. Perhaps unsurprisingly, randomization along with CD can provide constant throughput (in non-jammed slots) despite an unbounded adversary; even without CD, jamming can be tolerated, although throughput decreases slightly to $\Theta(1/\log n)$. When the adversary is bounded, a range of results exist. Under a saturated model, randomized results demonstrate constant throughput for a variety of network settings against a bounded adversary, so long as some weak information about the adversary’s power and $n$ is known. In contrast, deterministic results in the AQT setting are more nuanced, being parameterized by aspects of jamming rate and burstiness, and generally more pessimistic.
Existing results that tolerate jamming exhibit a sharp trade-off between the amount of listening devices perform and throughput achieved. On one hand, if each participant listens to the channel in every slot (except when sending), then constant throughput (in expectation or w.h.p. in non-jammed slots) is possible (bender:how, ; BenderFGY19, ; ChangJP19, ; awerbuch:jamming, ; richa:jamming2, ; richa:jamming3, ). Note that such results are not energy efficient, given that listening is roughly as costly as sending in practice (recall Section 1.1.3). On the other hand, if devices never listen to the channel, only sub-constant throughput is possible, and the best result is $O(1/\log n)$ throughput w.h.p. (DBLP:conf/podc/ChenJZ21, ). The recent result in (bender:fully, ) demonstrates that jamming-resistant contention resolution is possible with constant throughput and a polylogarithmic number of channel accesses (both sending and listening) without using control bits. However, it is unclear whether better energy-efficiency bounds are possible in the presence of jamming, perhaps under models with additional feedback permitted on the channel.
5. Conclusion
We have surveyed the literature on adversarial contention resolution by delineating results based on key model assumptions, metrics, and algorithm-design choices. The problem of contention resolution remains an active topic of research, and here we briefly discuss some potential directions for future research.
Machine learning (ML) may have an important role to play in the design of contention-resolution algorithms; certainly, the use of ML is being explored in the development of future wireless standards, such as 6G (kato:ten, ). As discussed in Section 3.3.2, Gilbert et al. (gilbert2021contention, ) demonstrate the impact of system-size predictions, which might be offered by a black-box ML module. A natural question is whether there are other useful predictions that could give improved performance, or at least performance that is parameterized by the accuracy of the predictions. This has been done recently for denial-of-service attacks in the traditional client-server model (chakraborty:bankrupting, ). Can predictions about whether a slot will be noisy (or jammed) yield improved robustness?
In a similar vein, incorporating capabilities enabled by new (but existing) wireless technology can yield novel results (e.g., coded networks (bender:coded, )) and deserves further exploration. For instance, the 5G New Radio (NR) ability to dynamically change the width of the frequency band being used can help mitigate congestion or interference on the channel, and allow for power savings (e.g., (abinader:impact, )). Defining a clean abstract model for these capabilities and studying their effects on fundamental problems like contention resolution could be a valuable research focus.
In many wireless systems, collisions have significant cost. For example, in WiFi networks, the entire packet must be sent before the sender learns of success or failure (due to a collision). Under coded networks, multiple packets may be sent simultaneously; therefore, a collision ruins what could otherwise be multiple successful transmissions in a slot. In mobile settings, failure to transmit a packet may mean incurring a significant delay before the sender is next within range of the intended receiver. Therefore, modeling collision cost is another angle that presents some interesting open problems in the area of contention resolution.
Mobility adds a layer of complexity to wireless communication because participants may be constantly changing position, affecting signal strength and interference. This is an aspect that has not received attention in the literature, and integrating mobility into contention resolution might open the door to interesting new problems. Additional costs due to distance traveled or energy expended to overcome attenuation are aspects that may be worthwhile considering.
Adopting different channel models can yield distinct upper and lower bounds compared to the classical model, as we have seen in the case of the SINR model (e.g., (fineman:contention, ; jurdzinski2015cost, )). The power sensitive model explored by De Marco and Kowalski (DBLP:conf/icdcs/MarcoK10, ) and by Tsybakov tsybakov:multiplicity) is another example of a interesting channel model. In the latter case, exploring via experiment the degree to which this model holds in practice may be interesting, as it seems to impart a non-trivial amount of power (e.g., we can count the number of users in $O(1)$ slots rather than ${\texttt{poly}({\log n})}$ slots). Even coarser-grained feedback —rather an exact value—for the number of senders could be useful. For example, in the classical channel, it is hard to tell if noise comes from many users competing for access or from jamming; however, (even rough) estimates could help clarify this.
Another dimension to consider is traffic priority. Consider a packet that contains voice-over-IP data, and another packet holding old data being backed up on a remote server. Which packet should get access to the shared channel first? Intuitively, the first packet is time sensitive and should have priority; if it is delayed too long, the packet contents are useless. This prioritization will become increasingly important to wireless networks, as a combination of limited bandwidth and a growing number of devices forces a differentiated treatment of traffic. While this aspect has been explored in (DBLP:conf/spaa/AgrawalBFGY20, ), there are several outstanding questions regarding robustness to adversarial jamming, heterogeneous packet sizes, and potential trade-offs with the amount of delay packets may suffer as a result.
Energy efficiency is likely to remain an important consideration in wireless systems. The exact limits of trading off between robustness to worst-case noise, energy efficiency, and latency still remain unknown. Can we obtain $o(\log n)$ number of channel-access attempts per packet despite jamming, while retaining asymptotically optimal latency? Related to this aspect, in some wireless standards, transmission power can be increased (and decreased), allowing devices to compensate for signal attenuation. This capability might be useful in overcoming jamming attacks, where good participants incur higher energy costs to transmit at higher power in order to overcome malicious noise. Is it possible to design algorithms in the spirit of (ChangJP19, ; BenderFGY19, ), where the sending costs are grow very slowly with the jamming costs? Alternatively, can the latency be asymptotically improved over existing jamming-resistant results?